# Discovering and exploiting active sensing motifs for estimation

### v1

@@ Discussion
Our observability framework quantifies the minimum error variance for the value of an initial state variable given subsequent measurements, i.e. an error bound on $\hat{x}_{i,k}$ given $\mathbf{Y}_{k:k+\omega}$. In state estimation, we are generally interested in building state estimates given preceding measurements, that is, we want to find $\hat{x}_{i,k}$ given $\mathbf{Y}_{k-\omega:k}$. To quantify the error bound for this final state, it would be more accurate to use one of observability’s dual principles, namely constructability Boyacoglu2024, also referred to as the posterior Cramér Rao bound tichavsky1998posterior, or observability of the final state sontag1990. We chose to analyze observability, rather than the mathematically dual concept of constructability (which analyzes the final state, rather than the initial state Boyacoglu2024) because it allowed us to use an initial state perturbation approach for empirically calculating the observability matrix $\mathcal{O}$. This perturbation approach is efficient, and works seamlessly with non-closed-form models. For the purpose of discovering active sensing motifs, both approaches will provide equivalent insight. In cases where a quantification of the error bounds for the final state is necessary, our calculations of $\mathcal{F}$ can be replaced with an iterative equation that uses a discretized time-varying approximation of the continuous-time dynamics Boyacoglu2024. Taking this approach would also allow for the inclusion of non-zero process noise covariance matrix $Q$ to understand how uncertainty of an agent’s internal model affects the observability (or constructability) of state variables of interest. We note that for the relatively short time windows and small model covariances ($Q\approx 0$) that we considered, the observability and constructability will generally be strongly correlated, allowing us to use our empirical observability calculations as a proxy for constructability. For our demonstration of the AI-KF we circumvented this issue by empirically establishing the error bound of a data-driven estimator by analyzing its error variance directly.
Limitations and future directions of the AI-KF
One limitation that will need to be addressed in the future is that the AI-KF may artificially shrink error covariances during periods of high observability, particularly if multiple data-driven estimators are used to augment the KF. This could be mitigated by only using augmented information when the AI-KF and UKF disagree, and by carefully considering the covariance of data-driven estimators that have coupled performance (i.e. off diagonals of $\mathcal{R^{\prime}}$). We note that the UKF also fails to provide accurate estimates of covariance, particularly during periods of low observability (Fig. LABEL:fig:figure_4_d), something that future work with observability or constructability metrics may also help to address.
Active sensing in organisms
Perhaps the most critical component to active sensing is knowing when to do it. Constantly performing active sensing maneuvers can be wasteful, or may conflict with other navigation goals. To achieve a balance, many animals appear to alternate between these two goals in an exploration vs. exploitation trade-off that is triggered when estimates reach a threshold of error Biswas2023. The observability-informed filter approach we introduced provides a flexible and parsimonious framework for taking advantage of observability information (or an approximation thereof) to intelligently update state estimates only during periods of intermittent active sensing (Figure 3). Whether animals use something resembling an observability filter or AI-KF remains an open question. However, Drosophila have been shown to throttle learning plasticity based on a turning signal Fisher2022a, a motif that we find generally corresponds to high levels of observability for many state variables (Fig. LABEL:fig:figure_2_c). In many regions of the brain multiple sensory signals are combined together, often with low- or high-pass filter dynamics, making it difficult to understand what role these signals may play. Our approach can be used to understand which task relevant state variables these multiplexed signals can estimate, thereby providing hypotheses about their role in navigation May2025compact.
Combining data-driven and model-based estimation in robotics
Model-based state estimation is the standard for autonomous systems urrea2021kalman. However, for challenging estimation tasks there are clear advantages to methods that incorporate time windows of sensor information through nonlinear optimization approaches (e.g. wind estimation VanBreugel2022 and monocular SLAM Shen2015), or more recent data-driven machine learning methods pellerito2024deep. Such methods, however, rarely confer advantages that traditional methods provide such as stability guarantees and computational simplicity. The AI-KF provides a bridge between these two worlds by outlining a statistically meaningful approach for blending model-based and data-driven estimation. The approach is inherently modular, making it easy for individual data-driven estimators to be added or removed from the AI-KF without requiring retraining or retuning. For example, existing monocular SLAM (Simultaneous Localization and Mapping) estimates Weiss2011, or deep visual odometry estimates pellerito2024deep can be incorporated into the AI-KF without further modification. We do not suggest that the AI-KF will universally improve upon traditional methods, however its benefits emerge for systems with sparse and dynamic observability properties that are difficult for traditional methods to handle. Together, our BOUNDS and AI-KF methods can help engineers develop machines that use active sensing to minimize the number of required sensors, plan trajectories, and design state estimators that work harmoniously to enable levels of robustness commonly observed in organisms.
Methods
Observability calculations
Empirical observability matrix
We constructed empirical observability matrices $\mathcal{O}$ as previously described Cellini2023. In brief, for a given state trajectory $\mathbf{X}$ we set a window size $\omega$ and initial state $\mathbf{x}_{0}$ in sliding windows (LABEL:fig:figure_1_a). Then we perturbed each individual initial state variable by a small amount $\varepsilon=$1\text{\times}{10}^{-5}\text{\,}$\$ in both directions $x_{i,0}^{\pm}=x_{i,0}\pm\varepsilon$. The resulting differences in the sensor measurements $\mathbf{Y}$ (LABEL:fig:figure_1_c) was collected and normalized by the perturbation amount $\Delta\mathbf{Y}/2\varepsilon$ and arranged as rows in $\mathcal{O}$ (LABEL:fig:figure_1_e). All angular measurement variables were unwrapped to ensure that $\Delta\mathbf{Y}$ did not have any discontinuities, a prerequisite for empirical observability analysis. We set a window size of $\omega=5$ for all simulations.
Fisher information and its inverse
For each $\mathcal{O}$ we computed the Fisher information matrix as $\mathcal{F}=\mathcal{O}^{\intercal}\mathcal{R}^{-1}\mathcal{O}$. We set the block diagonal noise covariance matrix $\mathcal{R}=0.1\cdot I_{(p\times\omega)\times(p\times\omega)}$ where $p$ is the number of measurements and $\omega$ is the sliding window size. This assumes the noise variance on each measurement is either $0.1\text{\,}\mathrm{r}\mathrm{a}\mathrm{d}$ for angular measurements, $0.1\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$ for airspeed and ground speed magnitude measurements, and $0.1\text{\,}{\mathrm{s}}^{-1}$ for optic flow measurements. The absolute noise properties of these measurements do not have a large impact on our results because larger or smaller values of $\mathcal{R}$ simply scale $\mathcal{F}$, however relative changes in $\mathcal{R}$ for different measurements would have an effect. We do not consider the process noise $Q$ in this manuscript, but with some modifications to our method (see Boyacoglu2024) this can be taken into account as well. Process noise will have a relatively small effect for a small window size, so this is a reasonable assumption.
Calculating the inverse of $\mathcal{F}$ is often not straightforward due to the presence of unobservable directions, which results in a matrix that is not invertible. A pseudoinverse can be computed, but this results in misleading (often small, when they should be very large) values within the diagonal elements of $\mathcal{F}^{-1}$ for unobservable state variables. Therefore, we chose to add a small amount of observability to each direction in $\mathcal{F}$ to ensure computing a traditional inverse was possible (Equation 3). We refer to this inverse as the “Chernoff inverse” chernoff1953locally, following the naming convention of prior literature on matrix inverses rao1972generalized. Note that the regularization does not have to be of the form $\lambda I_{n\times n}$. A more general regularization of $\lambda J$, where $J$ is any $n\times n$ positive definite matrix, is allowable. For simplicity, we chose $J$ to be the identity matrix. This method results in a cap on the upper bound of the error noise covariance matrix equal to $1/\lambda$. For all systems, $\lambda I_{n\times n}$ must be carefully selected to ensure that it is small enough not to distort the true observability properties, while still large enough to avoid numerical instabilities associated with computing a numerical inverse. We set $\lambda=$1\text{\times}{10}^{-6}\text{\,}$\$ for all calculations.
@@ Methods
$\mathbf{\dot{x}}=\begin{bmatrix}\dot{x}\\ \dot{y}\\ \dot{z}\\ \dot{v}_{x}\\ \dot{v}_{y}\\ \dot{v}_{z}\\ \dot{\phi}\\ \dot{\theta}\\ \dot{\psi}\\ \dot{w}\\ \dot{\zeta}\\ \end{bmatrix}=\begin{bmatrix}v_{x}\cos(\psi)-v_{y}\sin(\psi)\\ v_{x}\sin(\psi)+v_{y}\cos(\psi)\\ v_{z}\\ u_{z}k_{z}\cos(\phi)\sin(\theta)-Ca_{x}+v_{y}\dot{\psi}\\ -u_{z}k_{z}\sin(\phi)-Ca_{y}-v_{x}\dot{\psi}\\ -u_{z}k_{z}\cos(\phi)\cos(\theta-Cv_{z}+g)\\ u_{\phi}k_{\phi}\\ u_{\theta}k_{\theta}\\ u_{\psi}k_{\psi}\\ u_{w}\\ u_{\zeta}\\ \end{bmatrix}$. (16)
The model describes the velocity of the agent in the body-level North-East-Down (NED) coordinate system. $v_{x},v_{y}]$ define translational velocities in the $XY$ plane that rotate with the agents yaw heading $\psi$ and $v_{z}$ describes the global $Z$ velocity. Note that the the derivatives of $v_{x},v_{y}]$ include the derivative of the rotating reference frame. The body-level frame was chosen primarily because it simplifies the control of trajectories. The position of the agent $[x,y,z]$ was defined in the global frame to more easily visualize trajectories. In general, the choice of reference frame does not affect observability results and only serves to simplify some of the intermediate steps in our framework.
The model has four inputs for flight control. The acceleration in the body frame $u_{z}$ serves to counteract acceleration from gravity $g$ (note: we use the symbol $g$ to refer to ground speed apart from this line). By changing roll $\phi$ and pitch $\theta$, the agent can redirect $u_{z}$ and control its velocity vector. The inputs $u_{\phi},u_{\theta},u_{\psi}]$ are angular rates that allow the agent to control roll, pitch, and yaw. Most flying systems use a control allocationJohansen2013 approach to transform these basic inputs into control commands for individual actuators, like motor or tilter commands. However, this level of detail is unnecessary for the purpose of observability analysis and only serves to complicate the control of the agent. Thus, we do not model the actuation associated with these inputs. We augmented the state vector with motor calibration coefficient states for each input $k_{z},k_{\phi},k_{\theta},k_{\psi}]$, which were all set equal to one. These augmented states allowed us to consider cases where the agent did not know part of its input dynamics.
@@
$\begin{bmatrix}{a_{x}}\\ {a_{y}}\\ \end{bmatrix}=\begin{bmatrix}v_{x}-w\cos{(\psi-\zeta)}\\ v_{y}+w\sin{(\psi-\zeta)}\\ \end{bmatrix}$, (17)
@@
To test the sensitivity of our observability analysis to modeling choices, we also employed a model with more realistic 3D dynamics Beard2008 (Figure S2). This model is distinct from the kinematic model described above in that it uses forces and torques as inputs and considers physical parameters like mass $m$, inertia $I_{\bullet}$, and drag forces $C_{\bullet}$ about each axis. The states remained the same as in Equation 16, but we added states for angular velocity $\omega_{x},\omega_{y},\omega_{z}]$.
$\mathbf{\dot{x}}=\begin{bmatrix}\dot{x}\\ \dot{y}\\ \dot{z}\\ \dot{v}_{x}\\ \dot{v}_{y}\\ \dot{v}_{z}\\ \dot{\phi}\\ \dot{\theta}\\ \dot{\psi}\\ \dot{\omega}_{x}\\ \dot{\omega}_{y}\\ \dot{\omega}_{z}\\ \dot{w}\\ \dot{\zeta}\\ \end{bmatrix}=\begin{bmatrix}v_{x}\cos(\psi)-v_{y}\sin(\psi)\\ v_{x}\sin(\psi)+v_{y}\cos(\psi)\\ v_{z}\\ \frac{1}{m}(u_{z}\cos(\phi)\sin(\theta)-Ca_{x})+v_{y}\dot{\psi}\\ \frac{1}{m}(-u_{z}\sin(\phi)-Ca_{y})-v_{x}\dot{\psi}\\ \frac{1}{m}(-u_{z}\cos(\phi)\cos(\theta)-Cv_{z}+mg)\\ \omega_{x}+\tan(\theta)(\omega_{y}\sin(\phi)+\omega_{z}\cos(\phi))\\ \omega_{y}\cos(\phi)-\omega_{z}\sin(\phi)\\ \omega_{y}\frac{\sin(\phi)}{\cos(\theta)}-\omega_{z}\frac{\cos(\phi)}{\cos(\theta)}\\ \frac{u_{\phi}}{I_{x}}+\frac{I_{y}-I_{z}}{I_{x}}\omega_{y}\omega_{z}\\ \frac{u_{\theta}}{I_{y}}+\frac{I_{z}-I_{x}}{I_{y}}\omega_{x}\omega_{z}\\ \frac{u_{\psi}}{I_{z}}+\frac{I_{x}-I_{x}}{I_{z}}\omega_{x}\omega_{y}\\ \dot{w}\\ \dot{\zeta}\\ \end{bmatrix}$}. (18)
@@
translation damping ($C$) $0.1\text{\,}\mathrm{kg}\text{\,}{\mathrm{s}}^{-1}$
@@
$\mathbf{h}(\mathbf{x})=\begin{bmatrix}\psi\\ \beta\\ \gamma\\ \eta\\ g\\ a\\ r\\ q\end{bmatrix}=\begin{bmatrix}\psi\\ \operatorname{atan2}(v_{y},v_{x})\\ \operatorname{atan2}(a_{y},a_{x})\\ \operatorname{atan2}(\dot{v}_{y},\dot{v}_{x})\\ \sqrt{{v_{x}^{2}+v_{y}^{2}}}\\ \sqrt{{a_{x}^{2}+a_{y}^{2}}}\\ g/z\\ \sqrt{{(\dot{v}_{x}-v_{y}\dot{\psi})^{2}+(\dot{v}_{y}+v_{x}\dot{\psi})^{2}}}\end{bmatrix}$}. (19)
The $v_{y}\dot{\psi}$ and $v_{x}\dot{\psi}$ terms are present to cancel out the terms from 18 that describe the derivative of the rotating reference frame, as a body mounted accelerometer would not pick up these terms.
@@
We simulated our model in Python with a nonlinear differential equation solver Fiedler2023. Even though our model was defined with continuous-time dynamics, we simulated each time step discretely with constant inputs for each step. This was done as opposed to interpolating inputs between time steps, as interpolation between steps can cause issues reproducing simulations starting from an arbitrary point along a trajectory due to the way numerical solvers operate. Our discrete time approach ensured that we could start from any given point along a trajectory and maintain the exact simulation thereafter, which was necessary to construct observability matrices in sliding windows.
We used model predictive control (MPC) with full state feedback to precisely control flight trajectories through our model (Equation 18) (LABEL:fig:figure_1_b). In principle, other control methods could work, but MPC is a good general choice for controlling nonlinear systems. We first prescribed—or took from measured trajectories—a desired set-point for $v_{x}$, $v_{y}$, and $\psi$ over time (LABEL:fig:figure_1_a), then used an open-source MPC toolbox Fiedler2023 to find the optimal control inputs to track the set-point variables. Our MPC cost function penalized the squared error between each of our set-point series while adding an additional penalty for each control input to ensure more smoother force and torque inputs. Specifics can be found in our publicly available code.
@@
We designed a feed-forward artificial neural network (ANN) for regression to estimate wind direction LABEL:fig:figure_3_a. There were 12 input neurons consisting of the heading $\psi$, apparent airflow angle $\gamma$, and optic flow angle $\beta$ in the time window $\omega=4$. We used three hidden layers with 64 neurons and an output layer with a single neuron to estimate wind direction $\zeta$. All neurons were fully connected and used a rectified linear unit activation function, except for the output layer, which used a linear activation function. We also added a Gaussian noise layer after the input layer ($\mu=0$, $\sigma=0.01$ rad), which was only active during training, to help the ANN deal with noise in the measurements. We built and trained our network using TensorFlow Abadi2016 and Keras Chollet2015. We designed a custom loss function to deal with the circular nature of the output variable:
@@
where $v_{x,\omega}/v_{x,0}$ represents an acceleration as the ratio of initial and final $v_{x}$. This was used instead of $\dot{v}_{x}$ to ensure flight was always generally directed forward and there was not backwards flight for small starting $v_{x}$ values. We applied a 80-20 train-test split to the data. Each trajectory was $0.2\text{\,}\mathrm{s}$ (21 discrete points). We augmented the measurements from each trajectory with $\omega=4$ prior time steps, so only the 17 points with $\omega$ prior points available were used in training/testing. The training and testing data split into equally sized bins, once sorted by observability and once sorted randomly (LABEL:fig:figure_3_c). We trained and tested multiple ANN’s on these binned datasets (Figure S4). We trained each ANN for 10000 epochs with a batch size of 4096. Specific details can be found in our publicly available code.
@@
The altitude estimator ANN was designed similarly to the wind direction estimator ANN (LABEL:fig:figure_supp_altitude_ANN_a). The ANN had 40 input neurons consisting of the forward optic flow $r_{x}$ and forward acceleration $\dot{v}_{z}$ in the time window $\omega=20$. Like the wind direction estimator, We used three hidden layers with 64 neurons and an output layer with a single neuron to estimate altitude $\zeta$.
@@
We used the MATLAB implementation of an Unscented Kalman Filter (which uses the utilize square-root factorization of covariance matrices for improved numerical stability) Wan2006 for all Augmented Information Kalman Filter simulations. The Unscented transform parameters were set to $\alpha=10^{-3},\beta=1,\kappa=0$. The initial state and covariance estimates, process noise covariance, and measurement noise were varied (Figure 4). The baseline noise variance for the optic flow measurement was set to $10^{-3}$. Specifics can be found in our publicly available code.
@@
We collected data from a quadcopter (DJI Matrice 300 RTK) flying outdoors in an open, grassy park area with relatively flat terrain near Reno, NV (LABEL:fig:figure_5_a). Acceleration data was measured by the onboard accelerometer and ventral camera images were collected at 60 Hz by a ventrally mounted Raspberry Pi camera. Optic flow was computed offline using an implementation of Dense Inverse Search in OpenCV (Python) and low-pass filtered with a cutoff frequency of 1 Hz opencv_library. We scaled the optic flow with a constant scalar (approximately 0.057) to account for camera and lens properties such that optic flow was equivalent to the ratio of velocity over altitude. Ground truth data was measured by an onboard GPS. All flights were operated in velocity control mode, with manually specified accelerations/decelerations.

### v2

@@ 7. Reveal active sensing motifs by correlating state dynamics with observability
Our observability framework quantifies the minimum error variance for the value of an initial state variable given subsequent measurements, i.e. an error bound on $\hat{x}_{i,k}$ given $\mathbf{Y}_{k:k+\omega}$. In state estimation, we are generally interested in building state estimates given preceding measurements, that is, we want to find $\hat{x}_{i,k}$ given $\mathbf{Y}_{k-\omega:k}$. To quantify the error bound for this final state, it would be more accurate to use one of observability’s dual principles, namely constructability 8, also referred to as the posterior Cramér-Rao bound 55, or observability of the final state 51. We chose to analyze observability, rather than constructability, because it allowed us to use an initial state perturbation approach for empirically calculating the observability matrix $\mathcal{O}$. This perturbation approach is efficient, and works seamlessly with non-closed-form models. For the purpose of discovering active sensing motifs, both approaches will provide equivalent insight. In cases where a quantification of the error bounds for the final state is necessary, our calculations of $\mathcal{F}$ can be replaced with an iterative equation that uses a discretized time-varying approximation of the continuous-time dynamics 8. Taking this approach would also allow for the inclusion of non-zero process noise covariance matrix $Q$ to understand how uncertainty of an agent’s internal model affects the observability (or constructability) of state variables of interest. We note that for the relatively short time windows and small model covariances ($Q\approx 0$) that we considered, the observability and constructability will generally be strongly correlated, allowing us to use our empirical observability calculations as a proxy for constructability. For our demonstration of the AI-KF we circumvented this issue by empirically establishing the error bound of a data-driven estimator by analyzing its error variance directly.
@@
Hyper-parameters
BOUNDS has two primary hyper-parameters, 1) the regularization parameter $\lambda$ (Eq. 13), and 2) the window size $\omega$ over which to analyze the observability. It is best practice to choose $\lambda$ to be large enough to numerically invert the Fisher information matrix, but small enough to not distort the total information. In general, a good guideline is to choose $\lambda$ to be less than the smallest nonzero eigenvalue of the Fisher information matrix 28. Alternatively, one could symbolically compute the limit in Eq. 13, however, this can be computationally expensive and can yield infinite values corresponding to unobservable state variables. For $\omega$, we recommend choosing a window size roughly equal to the slowest motif of interest.
@@
$\mathbf{\dot{x}}=\begin{bmatrix}\dot{x}\\ \dot{y}\\ \dot{z}\\ \dot{v}_{x}\\ \dot{v}_{y}\\ \dot{v}_{z}\\ \dot{\phi}\\ \dot{\theta}\\ \dot{\psi}\\ \dot{w}\\ \dot{\zeta}\\ \end{bmatrix}=\begin{bmatrix}v_{x}\cos(\psi)-v_{y}\sin(\psi)\\ v_{x}\sin(\psi)+v_{y}\cos(\psi)\\ v_{z}\\ u_{z}k_{z}\cos(\phi)\sin(\theta)-Ca_{x}+v_{y}\dot{\psi}\\ -u_{z}k_{z}\sin(\phi)-Ca_{y}-v_{x}\dot{\psi}\\ -u_{z}k_{z}\cos(\phi)\cos(\theta)-Cv_{z}+\mathscr{g}\\ u_{\phi}k_{\phi}\\ u_{\theta}k_{\theta}\\ u_{\psi}k_{\psi}\\ u_{w}\\ u_{\zeta}\\ \end{bmatrix}$. (17)
The model describes the velocity of the agent in the body-level North-East-Down (NED) coordinate system. $v_{x},v_{y}]$ define translational velocities in the $XY$ plane that rotate with the agents yaw heading $\psi$ and $v_{z}$ describes the global $Z$ velocity. Note that the derivatives of $v_{x},v_{y}]$ include the derivative of the rotating reference frame. The body-level frame was chosen primarily because it simplifies the control of trajectories. The position of the agent $[x,y,z]$ was defined in the global frame to more easily visualize trajectories. In general, the choice of reference frame does not affect observability results and only serves to simplify some of the intermediate steps in our framework.
The model has four inputs for flight control. The acceleration in the body frame $u_{z}$ serves to counteract acceleration from gravity $\mathscr{g}$ (note: we use the symbol $g$ to refer to ground speed, and $\mathscr{g}$ for acceleration from gravity, and do not refer to gravity anywhere except this line and the $\dot{v}_{z}$ dynamics). By changing roll $\phi$ and pitch $\theta$, the agent can redirect $u_{z}$ and control its velocity vector. The inputs $u_{\phi},u_{\theta},u_{\psi}]$ are angular rates that allow the agent to control roll, pitch, and yaw. Most flying systems use a control allocation30 approach to transform these basic inputs into control commands for individual actuators, like motor or tilter commands. However, this level of detail is unnecessary for the purpose of observability analysis and only serves to complicate the control of the agent. Thus, we do not model the actuation associated with these inputs. We augmented the state vector with motor calibration coefficient states for each input $k_{z},k_{\phi},k_{\theta},k_{\psi}]$, which were all set equal to one. These augmented states allowed us to consider cases where the agent did not know part of its input dynamics.
@@
$\begin{bmatrix}{a_{x}}\\ {a_{y}}\\ \end{bmatrix}=\begin{bmatrix}-v_{x}+w\cos{(\psi-\zeta)}\\ -v_{y}-w\sin{(\psi-\zeta)}\\ \end{bmatrix}$, (18)
@@
To test the sensitivity of our observability analysis to modeling choices, we also employed a model with more realistic 3D dynamics 3 (Figure S2). This model is distinct from the kinematic model described above in that it uses forces and torques as inputs and considers physical parameters like mass $m$, inertia $I_{\bullet}$, and drag forces $C_{\bullet}$ about each axis. The states remained the same as in Equation 17, but we added state variables for angular velocity $\omega_{x},\omega_{y},\omega_{z}]$.
$\mathbf{\dot{x}}=\begin{bmatrix}\dot{x}\\ \dot{y}\\ \dot{z}\\ \dot{v}_{x}\\ \dot{v}_{y}\\ \dot{v}_{z}\\ \dot{\phi}\\ \dot{\theta}\\ \dot{\psi}\\ \dot{\omega}_{x}\\ \dot{\omega}_{y}\\ \dot{\omega}_{z}\\ \dot{w}\\ \dot{\zeta}\\ \end{bmatrix}=\begin{bmatrix}v_{x}\cos(\psi)-v_{y}\sin(\psi)\\ v_{x}\sin(\psi)+v_{y}\cos(\psi)\\ v_{z}\\ \frac{1}{m}(u_{z}\cos(\phi)\sin(\theta)-Ca_{x})+v_{y}\dot{\psi}\\ \frac{1}{m}(-u_{z}\sin(\phi)-Ca_{y})-v_{x}\dot{\psi}\\ \frac{1}{m}(-u_{z}\cos(\phi)\cos(\theta)-Cv_{z}+m\mathscr{g})\\ \omega_{x}+\tan(\theta)(\omega_{y}\sin(\phi)+\omega_{z}\cos(\phi))\\ \omega_{y}\cos(\phi)-\omega_{z}\sin(\phi)\\ \omega_{y}\frac{\sin(\phi)}{\cos(\theta)}-\omega_{z}\frac{\cos(\phi)}{\cos(\theta)}\\ \frac{u_{\phi}}{I_{x}}+\frac{I_{y}-I_{z}}{I_{x}}\omega_{y}\omega_{z}\\ \frac{u_{\theta}}{I_{y}}+\frac{I_{z}-I_{x}}{I_{y}}\omega_{x}\omega_{z}\\ \frac{u_{\psi}}{I_{z}}+\frac{I_{x}-I_{x}}{I_{z}}\omega_{x}\omega_{y}\\ \dot{w}\\ \dot{\zeta}\\ \end{bmatrix}$}. (19)
@@
translational damping ($C$) $0.1\text{\,}\mathrm{kg}\text{\,}{\mathrm{s}}^{-1}$
@@
$\mathbf{h}(\mathbf{x})=\begin{bmatrix}\psi\\ \beta\\ \gamma\\ \eta\\ g\\ a\\ r\\ q\end{bmatrix}=\begin{bmatrix}\psi\\ \operatorname{atan2}(v_{y},v_{x})\\ \operatorname{atan2}(a_{y},a_{x})-2\pi\\ \operatorname{atan2}(\dot{v}_{y},\dot{v}_{x})\\ \sqrt{{v_{x}^{2}+v_{y}^{2}}}\\ \sqrt{{a_{x}^{2}+a_{y}^{2}}}\\ g/z\\ \sqrt{{(\dot{v}_{x}-v_{y}\dot{\psi})^{2}+(\dot{v}_{y}+v_{x}\dot{\psi})^{2}}}\end{bmatrix}$}. (20)
The $v_{y}\dot{\psi}$ and $v_{x}\dot{\psi}$ terms are present to cancel out the terms from Equation 19 that describe the derivative of the rotating reference frame, as a body mounted accelerometer would not pick up these terms. The $2\pi$ term for $\gamma$ is necessary because $\gamma$ is defined in 2 with the vector pointed towards the agent, rather than away.
@@
We simulated our continuous-time dynamics models in Python with an ordinary differential equation (ODE) solver 21. As opposed to applying the ODE solver once across the full time-series, we applied it to each discrete time step independently while holding the inputs constant at each step. This was done as opposed to interpolating inputs between time steps, as interpolation between steps can cause issues reproducing simulations starting from an arbitrary point along a trajectory due to the way numerical solvers operate. Our discrete time approach ensured that we could start from any given point along a trajectory and maintain the exact simulation thereafter, which was necessary to construct observability matrices in sliding windows.
We used model predictive control (MPC) with full state feedback to precisely control flight trajectories through our model (Equation 19) (LABEL:fig:figure_1_b). In principle, other control methods could work, but MPC is a good general choice for controlling nonlinear systems. We first prescribed—or took from measured trajectories—a desired set-point for $v_{x}$, $v_{y}$, and $\psi$ over time (LABEL:fig:figure_1_a), then used an open-source MPC toolbox 21 to find the optimal control inputs to track the set-point variables. Our MPC cost function penalized the squared error between each of our set-point series while adding an additional penalty for each control input to ensure smoother force and torque inputs. Specifics can be found in our publicly available code.
@@
We designed a feed-forward artificial neural network (ANN) for regression to estimate wind direction LABEL:fig:figure_3_a. There were 12 input neurons consisting of the heading $\psi$, apparent airflow angle $\gamma$, and optic flow angle $\beta$ in the time window $\omega=4$. We used three hidden layers with 64 neurons and an output layer with a single neuron to estimate wind direction $\zeta$. All neurons were fully connected and used a rectified linear unit activation function, except for the output layer, which used a linear activation function. We also added a Gaussian noise layer after the input layer ($\mu=0$, $\sigma=0.01$ rad), which was only active during training, to help the ANN deal with noise in the measurements. We built and trained our network using TensorFlow 1 and Keras 40. We designed a custom loss function to deal with the circular nature of the output variable:
@@
where $v_{x,\omega}/v_{x,0}$ represents an acceleration as the ratio of initial and final $v_{x}$. This was used instead of $\dot{v}_{x}$ to ensure flight was always generally directed forward and there was not backwards flight for small starting $v_{x}$ values. We applied a 80-20 train-test split to the data. Each trajectory was $0.2\text{\,}\mathrm{s}$ (21 discrete points). We augmented the measurements from each trajectory with $\omega=4$ prior time steps, so only the 17 points with $\omega$ prior points available were used in training/testing. The training and testing data split into equally sized bins, once sorted by observability and once sorted randomly (LABEL:fig:figure_3_c). We trained and tested multiple ANNs on these binned datasets (Figure S4). We trained each ANN for 10000 epochs with a batch size of 4096. Specific details can be found in our publicly available code.
@@
The altitude estimator ANN was designed similarly to the wind direction estimator ANN (LABEL:fig:figure_supp_altitude_ANN_a). The ANN had 40 input neurons consisting of the forward optic flow $r_{x}$ and forward acceleration $\dot{v}_{x}$ in the time window $\omega=20$. Like the wind direction estimator, we used three hidden layers with 64 neurons and an output layer with a single neuron to estimate altitude $z$.
@@
Implementation details
We used the MATLAB implementation of an Unscented Kalman Filter (which uses the utilize square-root factorization of covariance matrices for improved numerical stability) 63 for all Augmented Information Kalman Filter simulations. The Unscented transform parameters were set to $\alpha=10^{-3},\beta=1,\kappa=0$. The initial state and covariance estimates, process noise covariance, and measurement noise were varied (Figure 4). The baseline noise variance for the optic flow measurement was set to $10^{-3}$. Specifics can be found in our publicly available code.
Eliminating double counting of information
After the AI-KF has converged, and during active sensing maneuvers yielding trustworthy (i.e. highly observable; low $\operatorname{Var}(\tilde{\check{x}}_{i})$) values for the data driven estimates, $\check{x}_{i}$ and $\hat{x}_{i-1}$ will not necessarily be independent. This breaks assumptions of the Kalman framework and would lead to an artificially low error variance for the estimates. This can be resolved through a small modification of $\mathcal{R}^{\prime}$ with the following steps.
  1. 1.
    Determine two estimates: a naive KF estimate $\hat{\text{\@text@baccent{$x$}}}_{i}$ that ignores the augmented terms (i.e. by setting $\operatorname{Var}(\tilde{\check{x}}_{i})$ to a large number approaching infinity, and the off-diagonal elements to zero), and the data-driven estimate $\check{x}_{i}$.
  2. 2.
    Calculate the difference between these two estimates:
    $d_{i}=\hat{\text{\@text@baccent{$x$}}}_{i}-\check{x}_{i}.$ (21)
  3. 3.
    Define a “relevance ratio”:
    $r_{i}=\frac{d_{i}^{2}}{\operatorname{Var}(\tilde{\check{x}}_{i})}.$ (22)
    When $r_{i}$ is small it means that the data-driven estimate carries little unique information, and should therefore be ignored. When $r_{i}$ is large, the data-driven estimate carries new information and should be used to augment the KF.
  4. 4.
    Define a throttling function that can be used to suppress the augmented measurement,
    $g(r_{i})=1+\frac{c}{r_{i}+\epsilon},\quad c>0,\quad\epsilon>0,$ (23)
    where $c$ is a tunable hyperparameter that controls the strength of suppression, and $\epsilon$ prevents division by zero and sets a ceiling on $g(r_{i})$. When the data-driven estimate carries no new information $g(r_{i})$ is very large (approaching $c/\epsilon$). When the data-driven estimate does carry new information $g(r_{i})$ approaches $1$.
  5. 5.
    Multiply the $\operatorname{Var}(\tilde{\check{x}}_{i})$ in $\mathcal{R}^{\prime}$ by $g(r_{i})$, and divide the off-diagonal elements by $\mathcal{R}^{\prime}$ by $g(r_{i})$.
This framework ensures that when the data-driven estimate $\check{x}_{i}$ is both accurate (as indicated by $\operatorname{Var}(\tilde{\check{x}}_{i})$) and different from the naive KF estimate (i.e. $|d_{i}|\gg 0$, the AI-KF takes full advantage of the augmented information. However, when the KF has converged and produces estimates that are statistically indistinguishable to the data-driven estimates, the AI-KF automatically relaxes to a naive KF, eliminating the possibility of double counting information.
@@
We collected data from a quadcopter (DJI Matrice 300 RTK) flying outdoors over relatively flat terrain near Reno, NV (LABEL:fig:figure_5_a). Acceleration data was measured by the onboard accelerometer and ventral camera images were collected at 60 Hz by a ventrally mounted Raspberry Pi camera. Optic flow was computed offline using an implementation of Dense Inverse Search in OpenCV (Python) and low-pass filtered with a cutoff frequency of 1 Hz 9. We scaled the optic flow with a constant scalar (approximately 0.057) to account for camera and lens properties such that optic flow was equivalent to the ratio of velocity over altitude. Ground truth data was measured by an onboard GPS. All flights were operated in velocity control mode, with manually specified accelerations/decelerations.